If f(x) = 5x - 7, what is f^-1(x)?
Answer by jenrobrody(19)

If f(x) = 5x - 7, what is f^-1(x)?
f(x)=5x-7 can also be written as:
y=5x-7, now interchange x and y, and solve for y:
x=5y-7
+7 +7
x+7=5y
y=x/5+7/5
so f^-1(x)=x/5+7/5